A Carnot engine is operated between 2 reservoirs at temperature 177°C and 77°C.  If the engine receives 4200 Jules of Heat energy from the source in each cycle, calculate the amount of heat rejected to the sink in each cycle.  Calculate the efficiency and work done by the engine.

Grade:11

1 Answers

ANGAPALI SARKAR
5 years ago
here, T1 =(177+273)=450K and T2 = (77+273)=350 K. Efficiency is 1-(T2/T1)
                                                                                                          =1-(350/450)
                                                                                                          =2/9
Again,we know, efficiency =(work/heat supplied)
Hence, 2/9 =W/4200
     or , W =8400/9 = 933.33 J